The perimeter of a rectangle is 50cm. The length of the rectangle is one more than 3 times the width of the rectangle. What are the dimensions of the rectangle
Answer by Greenfinch(383)

Call the width W, then the length is 3W +1
Perimeter is thus 2(W + 3W + 1) = 50
4W + 1 = 25
4W = 24
W = 6 cm Length is 3W + 1 = 19 cm